# Bridging Domain Gap for Flight-Ready Spaceborne Vision

### v1

@@ 5.3.6 Data Augmentation
The experiments so far used all three data augmentation techniques—Style Augmentation, DeepAugment and RandConv. In order to verify the individual contribution of these augmentations, Table 7 reports the mean orientation errors for different data augmentation configurations. The default baseline only consists of 5 random augmentations from the Albumentations library implemented in the style of RandAugment [105] (section 4.3).
@@ 5.4 Summary
The previous section investigates different aspects of designing and training a robust pose estimation NN. Under the same training conditions with identical input image resolutions, EfficientDet (EfficientNet) consistently outperforms all other models in terms of the OOD robustness as measured on the SPEED+ HIL domains. On the other hand, ViTPose (ViT) dominates the training and inference latency for different model sizes but with inferior OOD robustness. However, various ablation studies reveal that the ViTPose models can be made as robust as EfficientDet with improved pre-training of the ViT backbone, extensive data augmentation and increased input image resolution. Specifically, even with twice higher image resolution, the inference of ViTPose (ViT-S/16) is still faster than that of EfficientDet (EfficientNet-B0) that is 1/5 in size. Doubling the input resolution does lead to the increase of training latency by a factor of 4, but as stated in Requirement #1 (section 3.1), the increased training time is less likely to become as mission-critical as the inference time on a satellite avionic with an onboard GPU. Moreover, ViTPose satisfies Requirement #3 by design (section 3.3), completely lacking BN layers that could interfere with accurate online training on single images.
Moving forward, the ViTPose models trained with the data augmentation proposed in section 4.3 are referred to as SPNv3 for convenience. The default SPNv3 configuration is the ViT backbone pre-trained via DeiT III [124] on IN-22K then fine-tuned on IN-1K, patch size $P=16$, and input resolution 448 $\times$ 448.
@@ 5.4.1 Comparison on SPEED+
Table 8 compares the final performances of SPNv3-S, a mid-size variant that takes the ViT-S/16 backbone (22.7M parameters), against the top entries of SPEC2021 [17] and other state-of-the-art models introduced during the post-mortem competition following SPEC2021 and others. It can be seen that, with a higher input resolution at 448 $\times$ 448, SPNv3-S already outperforms the top entries of the lightbox category. Moreover, ensembling the output heatmap predictions from three models trained with different random seeds further improves the overall performance, achieving $E_{\text{q}}=2.69^{\circ}$ on lightbox and $E_{\text{q}}=3.89^{\circ}$ on sunlamp. Some of the pose predictions by the ensemble are visualized in Fig. 7 via reprojection of the Tango wireframe model. Note that the winning teams for respective HIL categories—TangoUnchained and lava1302— adopted adversarial training, directly utilizing the unlabeled images of the HIL domains. On the other hand, the proposed SPNv3 does not utilize any information about the HIL domain during the training phase. The performance of SPNv3 is comparable to that of EagerNet [80] which also does not involve the HIL domain images during training. Note that EagerNet uses a pre-trained ConvNeXt-Base network with 89M parameters and generates multiple pose hypotheses from dense predictions of model coordinates and predicted errors, which are further refined iteratively using a region-based approach. On the other hand, SPNv3-S is smaller and simpler, predicting keypoint locations from lower-resolution heatmaps after just a single forward propagation.
In the end, the ensemble performance with 448 $\times$ 448 inputs would rank SPNv3-S at first place in the lightbox category and third place in the sunlamp category of SPEC2021 [17]. Indeed, running an ensemble of multiple models may be computationally expensive especially onboard the satellite avionics. However, considering that the inference of SPNv3-S with 448 $\times$ 448 inputs only takes about 34 ms on a GPU of an NVIDIA Jetson Nano (see Table 6), it may be possible to run an ensemble of a few models in real-time during RPOD. The largest variant of equivalent size, SPNv3-B with 86.3M parameters, marginally improves the accuracy, nearly matching that of EagerNet on lightbox yet still remaining in third place on sunlamp. However, such a difference in performance between SPNv3-S and SPNv3-B would likely not yield significant improvement that justifies a near quadruple increase in size when processed by an onboard navigation filter.
@@ 5.5 Limitations
The proposed SPNv3 based on ViTPose is not without limitations. One major limitation is its dependence on a pre-trained ViT backbone. More specifically, Table 5 reveals that the OOD robustness of the downstream pose estimation task depends not just on what dataset the ViT backbone is pre-trained on but also on how the pre-training was conducted. This observation implies that those who cannot afford the computational power nor the expertise of ImageNet pre-training are restricted in exploring and fine-tuning different architectural options. In fact, the dependence on ImageNet pre-training prevents this work from exploring other aspects of ViT backbones, e.g., varying the patch sizes ($P$). Furthermore, the optimally pre-trained backbone may not exist, in which case the sub-optimally pre-trained alternative may have to be used for training on SPEED+. In response to such cases, the authors’ previous work [19] shows how one can further train the sub-optimal NN onboard the satellite avionics in real-time during RPOD. However, if a NN can be trained according to the recipe analyzed in this section to be as OOD robust as possible, the online training on satellite avionics may be minimized or skipped altogether.
6 Conclusion
@@
This paper presented SPNv3, a computationally efficient and robust Neural Network (NN) model based on the Vision Transformer (ViT) architecture. SPNv3 is trained exclusively on computer-generated synthetic images to achieve state-of-the-art robustness on unknown spaceborne imagery. Then, it is tested exhaustively on the hardware-in-the-loop images from a robotic facility that can simulate high-fidelity spaceborne illumination conditions. Extensive analyses are performed to identify data augmentation, transfer learning, NN size and input image resolution as key aspects of the design and training that contribute the most to its Ouf-Of-Distribution (OOD) robustness. In the end, a mid-size SPNv3-S with 22.7M parameters can perform inference on a larger 448 $\times$ 448 input images at nearly 30Hz on a restrictive GPU environment, which is well above the nominal update frequency of modern satellite guidance, navigation and control systems. Therefore, this paper successfully demonstrated that SPNv3 is flight-ready, capable of achieving state-of-the-art robustness while simultaneously meeting the prescribed computational requirements. Moreover, additional experiments reveal that SPNv3 can be trained to operate at a distance as far as where the target spacecraft occupies merely 10 $\times$ 10 pixels region albeit with more outliers.
@@
The immediate next step is to address the requirement of power consumption by minimizing the memory footprint of NN while maintaining the OOD robustness of larger SPNv3 variants. Candidate methods include NN weight pruning and quantization to reduced-precision formats. Furthermore, the OOD robustness evaluation in this work relies on the SPEED+ dataset which is limited to a 10 m separation between the camera and the target due to the hardware restrictions. Therefore, the OOD robustness must be evaluated at a farther distance which predicates on improved calibration of the robotic testbed.
@@
This work is supported by the US Space Force SpaceWERX Orbital Prime Small Business Technology Transfer (STTR) contract number FA8649-23-P-0560 awarded to TenOne Aerospace in collaboration with SLAB.

### v2

@@ 6.3.6 Data Augmentation
The experiments so far used all three data augmentation techniques—Style Augmentation, DeepAugment and RandConv. In order to verify the individual contribution of these augmentations, Table˜7 reports the mean orientation errors for different data augmentation configurations. The default baseline only consists of 5 random augmentations from the Albumentations library implemented in the style of RandAugment [111] (section˜5.2).
@@ 6.4 Summary
The previous section investigates different aspects of designing and training a robust pose estimation NN. Under the same training conditions with identical input image resolutions, EfficientDet (EfficientNet) consistently outperforms all other models in terms of the OOD robustness as measured on the SPEED+ HIL domains. On the other hand, ViTPose (ViT) dominates the training and inference latency for different model sizes but with inferior OOD robustness. However, various ablation studies reveal that the ViTPose models can be made as robust as EfficientDet with improved pre-training of the ViT backbone, extensive data augmentation and increased input image resolution. Specifically, even with twice higher image resolution, the inference of ViTPose (ViT-S/16) is still faster than that of EfficientDet (EfficientNet-B0) that is 1/5 in size. Doubling the input resolution does lead to the increase of training latency by a factor of 4, but as stated in Requirement #1 (section˜3.1), the increased training time is less likely to become as mission-critical as the inference time on a satellite with an onboard GPU. Moreover, ViTPose satisfies Requirement #3 by design (section˜3.3), completely lacking BN layers that could interfere with accurate online training on single images.
Moving forward, the ViTPose models trained with the data augmentation proposed in section˜5.2 are referred to as SPNv3 for convenience. The default SPNv3 configuration is the ViT backbone pre-trained via DeiT III [130] on IN-22K then fine-tuned on IN-1K, patch size $P=16$, and input resolution 448 $\times$ 448.
@@ 6.4.1 Comparison on SPEED+
Table˜8 compares the final performances of SPNv3-S, a small variant that takes the ViT-S/16 backbone (22.7M parameters), against the top entries of SPEC2021 [21] and other state-of-the-art models introduced during the post-mortem competition following SPEC2021 and others. It can be seen that, with a higher input resolution at 448 $\times$ 448, SPNv3-S already outperforms the top entries of the lightbox category. Moreover, ensembling the output heatmap predictions from three models trained with different random seeds further improves the overall performance, achieving $E_{\text{q}}=2.69^{\circ}$ on lightbox and $E_{\text{q}}=3.89^{\circ}$ on sunlamp. Some of the pose predictions by the ensemble are visualized in Fig.˜8 via reprojection of the Tango wireframe model. Note that the winning teams for respective HIL categories—TangoUnchained and lava1302—adopted adversarial training, directly utilizing the unlabeled images of the HIL domains. On the other hand, the proposed SPNv3 does not utilize any information about the HIL domain during the training phase. The performance of SPNv3-S is also comparable to those of the post-mortem entries, surpassing FA-VAE [137] with 41.6M parameters (approximated from its DarkNet-53 backbone [138]) in both categories and Chen et al. [139], which adopts semi-supervised domain adaptation, in the lightbox category. The performances of different SPNv3 variants are also comparable to that of EagerNet [86], which does not involve the HIL domain images during training. Note that EagerNet uses a pre-trained ConvNeXt-Base [140] network with 89M parameters and generates multiple pose hypotheses from dense predictions of model coordinates and predicted errors, which are further refined iteratively using a region-based approach. On the other hand, SPNv3-S is smaller and simpler, predicting keypoint locations from lower-resolution heatmaps after just a single forward propagation.
@@
In the end, the ensemble performance with 448 $\times$ 448 inputs would rank SPNv3-S at first place in the lightbox category and third place in the sunlamp category of SPEC2021 [21]. Indeed, running an ensemble of multiple models may be computationally expensive especially onboard the satellite avionics. However, considering that the inference of SPNv3-S with 448 $\times$ 448 inputs only takes about 34 ms on a GPU of an NVIDIA Jetson Nano (see Table˜6), it may be possible to run an ensemble of a few models in real-time during RPOD. The largest variant of equivalent size, SPNv3-B with 86.3M parameters, marginally improves the accuracy, nearly matching that of EagerNet on lightbox yet still remaining in third place on sunlamp. However, such a difference in performance between SPNv3-S and SPNv3-B would likely not yield significant improvement that justifies a near quadruple increase in size when processed by an onboard navigation filter.
Finally, while this article focuses on the OOD robustness and onboard computational capability of SPNv3 as an independent module, the integrated performance of SPNv3 and AUKF, both with and without online supervised training, is available for interested readers in Park [141].
@@ 6.5 Limitations
The proposed SPNv3 based on ViTPose is not without limitations. One major limitation is its dependence on a pre-trained ViT backbone. More specifically, Table˜5 reveals that the OOD robustness of the downstream pose estimation task depends not just on what dataset the ViT backbone is pre-trained on but also on how the pre-training was conducted. This observation implies that those who cannot afford the computational power nor the expertise of ImageNet pre-training are restricted in exploring and fine-tuning different architectural options. In fact, the dependence on ImageNet pre-training prevents this work from exploring other aspects of ViT backbones, e.g., varying the patch sizes ($P$). Furthermore, the optimally pre-trained backbone may not exist, in which case the sub-optimally pre-trained alternative may have to be used for training on SPEED+. In response to such cases, the authors’ previous work [23] shows how one can further train the sub-optimal NN onboard the satellite avionics in real-time during RPOD. However, if a NN can be trained according to the recipe analyzed in this section to be as OOD robust as possible, the online training on satellite avionics may be minimized or skipped altogether.
7 Conclusions
@@
This paper presented SPNv3, a computationally efficient and robust NN model based on the ViT architecture. SPNv3 is trained exclusively on computer-generated synthetic images to achieve state-of-the-art robustness on unknown spaceborne imagery. Then, it is tested exhaustively on the hardware-in-the-loop images from a robotic facility that can simulate high-fidelity spaceborne illumination conditions. Extensive analyses are performed to identify data augmentation, transfer learning, NN size and input image resolution as key aspects of the design and training that contribute the most to its OOD robustness. In the end, a small SPNv3-S with 22.7M parameters can perform inference on larger 448 $\times$ 448 input images at nearly 30Hz on a restrictive GPU, which is well above the nominal update frequency of modern satellite guidance, navigation and control systems. Therefore, this paper successfully demonstrated that SPNv3 is flight ready, capable of achieving state-of-the-art robustness while simultaneously meeting the prescribed computational requirements.
@@
The immediate next step is to address the requirement of power consumption by minimizing the memory footprint of the NN while maintaining the OOD robustness of larger SPNv3 variants. Candidate methods include NN weight pruning and quantization to reduced-precision formats. Furthermore, the OOD robustness evaluation in this work relies on the SPEED+ dataset, which is limited to a 10 m separation between the camera and the target due to the hardware restrictions. Therefore, the OOD robustness must be evaluated at a farther distance which predicates on improved calibration of the robotic testbed.
@@
This work is supported by the US Space Force SpaceWERX Orbital Prime Small Business Technology Transfer (STTR) contract number FA8649-23-P-0560 awarded to TenOne Aerospace in collaboration with SLAB. The authors would like to thank OHB Sweden for the PRISMA images used in this work. The authors also thank Dr. Justin Kruger at Stanford’s Space Rendezvous Laboratory for providing an image of angles-only navigation from the NASA Starling mission.